Let f(x) = 3x. Find f( 3) and f(1/3).
Answer by jim_thompson5910(35256)

f(x) = 3x

f(3) = 3(3)

f(3) = 9

-------------------------
f(x) = 3x

f(1/3) = 3(1/3)

f(1/3) = 3/3

f(1/3) = 1

-------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim

-------------------------------------------------------------------------------------------------